Tatii:

What's (-12x+ 8)+(3x+19)??

Vocaloid:

since you have a + sign in front of the parentheses, you can drop the parentheses -12x + 8 + 3x + 19 combine like terms

Chickentender37:

(-12x+ 8)+(3x+19) remove the parenthesis -12x + 8 + 3x + 19 then simply add the like terms ( like terms are terms that have the same variable. an example would be 4a and 2a cause they both have a variable of a ) the like terms in -12x + 8 + 3x + 19 are -12x with 3x and 8 with 9 so simply add them and you will get the answer -12x + 3x = 8 + 9 =